Circular No. 7325

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

V1494 AQUILAE D. Moro, A. Pizzella, and U. Munari, Asiago Astrophysical Observa-

tory, write: “Spectra (1 nm/pixel; range 340–950 nm) of V1494 Aql (= N Aql 1999 No. 2) were obtained on Dec. 2.71 UT with the Asiago 1.22-m telescope (+ Boller & Chivens CCD spectrograph). The Balmer hydrogen series is in emission with a P-Cyg profile. The Hα emission component has a 1300 km/s FWHM; the absorption component is centered at −1850 km/s from the emission component. Fe II multiplet 42 at 492, 502, and 517 nm, and O I 777-nm, are visible in emission also with a P-Cyg profile.”

J. A. Mattei, AAVSO, reports an independent visual discovery of this nova by Gary T. Nowak, Essex Junction, VT, with 7×35 binoculars on Dec. 1.958 UT at mag 5.6. Additional m_{v} estimates: Dec. 2.68, 4.2 (L. Kiss, Szeged, Hungary); 2.803, 4.1 (G. M. Hurst, Basingstoke, England); 2.814, 4.2 (P. Schmeer, Bischmisheim, Germany); 2.921, 3.7 (J. G. de S. Aguiar, Campinas, Brazil); 2.940, 4.3 (D. W. E. Green, Cambridge, MA); 2.956, 3.9 (R. Lourencon, Jundiai, Brazil); 2.96, 4.1 (G. Seronik, Belmont, MA).

V382 VELORUM M. Orio, Turin Observatory and University of Wisconsin; A. N. Par-

mar, European Space Agency; M. Capalbi, Italian Space Agency; and L. Piro and T. Mineo, Consiglio Nazionale delle Ricerche, report on behalf of the BeppoSAX team: “BeppoSAX observed V382 Vel (N Vel 1999) during Nov. 20.180–20.974 UT. The nova has become a supersoft luminous x-ray source. The count rate was 0.097 ± 0.002 counts/s in the two MECS instruments (1.6–10 keV) and 3.469 ± 0.002 counts/s in the LECS instrument (0.1–10 keV). The spectrum is dominated by a blackbody-like component at ∼ 40 eV with unabsorbed flux ∼ 1.5–2 × 10^{−7} erg cm^{−2} s^{−1}, assuming a blackbody model and absorption around 2 × 10^{21} atoms/cm^{2}. At higher energy (1.0–10 keV), there is a harder component that can be approximately modeled by emission from a thermal plasma at kT ∼ 1 keV and flux a few × 10^{−12} erg cm^{−2} s^{−1}. At a distance of 2000 pc (Della Valle et al. 1999, IAUC 7193), the total bolometric luminosity is ∼ 10^{38} erg cm^{−2}. The x-ray spectrum has evolved dramatically since June (Orio et al., IAUC 7196; Mukai and Ishida, IAUC 7205), and the column of neutral hydrogen seems to have decreased by a factor of 10, due to thinning of the ejecta. Further observations at x-ray and other wavelength bands are encouraged.”

1999 December 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green